The amount of heat needed to raise the temperature of 5 moles of a monatomic gas from o degree to 50 degree keeping it's pressur constant is 
 

Grade:12th pass

1 Answers

Vikas Amritiya
2 years ago
At constant pressure,
Q =nCp​ΔT
207 =Cp​×10
Cp​=20.7
At constant volume, W = 0
Q = U =nCv​ΔT
Cp​=Cv​+R
Cv​=20.7−8.3=12.4
Q =12.4×10
Q = 124 J